QUESTION 2 The result for the following expression (Ⓡe-3t 8(t-2)dt = ? is C5 Ce-12 C2 Ce+12 C none of the above

JUST THE ANSWER
Answer #1

Formula f(t){t-a)dt = f(a). L2=**)(t-2)dt = e =*3)* = 2 ->